# Linear Discriminant Regularized Regression

### v1

@@ 3 Theoretical guarantees
As we mentioned earlier, we decided to focus on the case where the columns of $B$ are (potentially) sparse and allow for different sparsity patterns across its columns. Therefore, for some positive integer $s\leq p$, we consider the following parameter space of $B$:
@@
For ease of presentation, we simply use the largest number of non-zero entries among all columns of $B$. However, our analysis can be easily extended to $\|B_{\ell}\|_{0}\leq s_{\ell}$ by allowing different $s_{\ell}$ for $\ell\in[L]$. Moreover, the analysis of $\widehat{B}$ relies on the standard Restricted Eigenvalue condition (RE) on the matrix $\widehat{\Sigma}=n^{-1}{\mathbb{X}}^{\top}{\mathbb{X}}$. Since ${\mathbb{X}}$ is a matrix with random entries, we resort to the following quantity on the within-class covariance matrix $\Sigma_{w}$.
Definition 1.
For any integer $1\leq s\leq p$, define
Our analysis needs the following regularity conditions on $\Sigma_{w}$ and the matrix ${\mathbb{M}}$ of conditional means. For any matrix $M$, write $\|M\|_{\infty}=\max_{i,j}|M_{ij}|$.
@@
Assumption 3.
There exists an absolute constant $C<\infty$ such that $\|{\mathbb{M}}\|_{\infty}\leq C$.
A condition akin to 2 is commonly used in sparse linear regression models, see, for instance, Bühlmann and Van de Geer (2011); Giraud (2021). 3 is a mild regularity condition on the largest entry of ${\mathbb{M}}$. We remark that the requirements that $\kappa_{s}$, $\|\Sigma_{w}\|_{\infty}$ and $\|{\mathbb{M}}\|_{\infty}$ are absolutely bounded are not essential, and can be relaxed to allow dependence on $n$.
3.1 Finite sample bounds of the proposed regression matrix estimation
The following theorem states finite sample upper bounds of both the (in-sample and out-of-sample) prediction error and the estimation error of the regression matrix estimator $\widehat{B}$ in (1.14) with the lasso-penalty in (2.7). For any numbers $a,b\in\mathbb{R}$, we write $a\vee b=\max\{a,b\}$. For any sequences $a_{n}$ and $b_{n}$, we write $a_{n}\lesssim b_{n}$ if there exists some constant $C>0$ such that $a_{n}\leq Cb_{n}$, and $a_{n}\asymp b_{n}$ if $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$.
@@ 3.1 Finite sample bounds of the proposed regression matrix estimation
Under model (1.1) and (1.3) with Assumptions 13, assume that there exists some sufficiently small constant $c>0$ such that $B\in\Theta_{s}$ for some $1\leq s\leq p$ with
$(s\vee L)\log(n\vee p)\leq c~{}n$
Then, with probability at least $1-n^{-1}$, we have
$\max_{\ell\in[L]}{1\over\sqrt{n}}\|{\mathbb{X}}(\widehat{B}_{\ell}-B_{\ell})\|_{2}~{}\lesssim~{}\kappa_{s}^{-1/2}\lambda\sqrt{s}$
$\max_{\ell\in[L]}\|\Sigma^{1/2}(\widehat{B}_{\ell}-B_{\ell})\|_{2}~{}\lesssim~{}\kappa_{s}^{-1/2}\lambda\sqrt{s}$
$\max_{\ell\in[L]}\|\widehat{B}_{\ell}-B_{\ell}\|_{2}~{}\lesssim~{}\kappa_{s}^{-1}\lambda\sqrt{s}$
for the estimator $\widehat{B}$ obtained from (1.14) with any $\lambda$ satisfying
$\lambda\geq C\sqrt{\|\Sigma_{w}\|_{\infty}+\|{\mathbb{M}}\|_{\infty}^{2}}\sqrt{\log(n\vee p)\over nL}$ (3.2)
and some finite, large enough constant $C>0$.
@@
See Section A.2. ∎
The two main difficulties in our proofs are (a) to control $\max_{j\in[p]}\|{\bm{e}}_{j}^{\top}{\mathbb{X}}^{\top}({\mathbb{Y}}-{\mathbb{X}}B)\|_{\infty}$, and (b) to bound from below the restricted eigenvalue of the Gram matrix $\widehat{\Sigma}=n^{-1}{\mathbb{X}}^{\top}{\mathbb{X}}$. For (a), as pointed out in Gaynanova (2020), the difficulty in this model is elevated relative to the existing analysis in sparse linear regression models due to the fact ${\mathbb{Y}}$ is not a linear model in ${\mathbb{X}}$. We establish a sharp control of $\max_{j\in[p]}\|{\bm{e}}_{j}^{\top}{\mathbb{X}}^{\top}({\mathbb{Y}}-{\mathbb{X}}B)\|_{\infty}$ in Lemma 6. For (b), the existing result in Rudelson and Zhou (2012) is not readily applicable as rows of ${\mathbb{X}}$ are generated from a mixture of Gaussians. In Lemma 7 we establish a uniform bound of $v^{\top}(\widehat{\Sigma}-\Sigma)v$ over $s$-sparse vectors $v$ which, in conjunction with the reduction arguments in Rudelson and Zhou (2012), proves (b). Our analysis in this paper allows the number of classes $L$ to grow with the sample size $n$
Gaynanova (2020) considers a group-lasso regularized regression similar to (1.14) and derives the bound of the $\ell_{2}$-norm $\max_{j\in[p]}\|{\bm{e}}_{j}^{\top}{\mathbb{X}}^{\top}({\mathbb{Y}}^{\prime}-{\mathbb{X}}B^{\prime})\|_{2}$ for some different response matrix ${\mathbb{Y}}^{\prime}$ and different target matrix $B^{\prime}$. Their proof is different and in particular requires that $L$ is fixed, independent of $n$.
@@
3.2 Estimation of the discriminant directional matrix
We analyze the estimation of the discriminant matrix $B^{*}$ in this section. Recall that
$\widehat{B}^{*}=\widehat{B}\widehat{H}^{+}$
with $\widehat{H}$ obtained from (2.9). The following theorem states upper bounds of the transformed $(\widehat{B}^{*}-B^{*})$. Since $H^{-1}$ depends on ${\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}$ (see Lemma 9 of Appendix B for its proof), our results depend in turn on the quantity
$\Delta_{\infty}=\max_{\ell\in[L]}\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}.$
This quantity essentially reflects the largest Mahalanobis distance in the feature space between any two label classes as can be seen from the relation (under (1.2))
@@ 3.2 Estimation of the discriminant directional matrix
Under the conditions in Theorem 2, assume that $\Delta_{\infty}\geq 1$ and assume there exists some sufficiently small constant $c>0$ such that
$\Delta_{\infty}^{2}\sqrt{sL\log(n\vee p)\over n}\leq c.$ (3.3)
Then, with probability at least $1-n^{-1}$, we have
$\max_{\ell\in[L]}~{}\bigl{\|}\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\bigr{\|}_{2}\!\!~{}\lesssim~{}\Delta_{\infty}^{5/2}\sqrt{sL\log(n\vee p)\over n}.$
@@
See Section A.3. ∎
Remark 3.
Condition (3.3) ensures that, with high probability, $\lambda_{K}(\widehat{H})\geq\lambda_{K}(H)/2$ holds, and hence $\widehat{H}$ is invertible. In Theorem 4 below we will see that condition (3.3) is also needed for consistent estimation of the discriminant function. The reasonable condition $\Delta_{\infty}\geq 1$ requires that the largest separation among all pairs of conditional distributions does not vanish. It is assumed to simplify the presentation, and our analysis can be easily extended to the case $\Delta_{\infty}\to 0$ as $n\to\infty$.
3.3 Finite sample bounds of the misclassification rates
We first analyze the estimation of the discriminant functions in (2.1). For any $X$ that is distributed according to (1.1), we have
$\widehat{G}_{\ell}(X)-G_{\ell}(X)$ $=(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}_{\ell}+(\widehat{\mu}_{\ell}-2X)^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})+2\log\left({\pi_{\ell}/\widehat{\pi}_{\ell}}\right).$
As a consequence of Theorem 3, the Gaussian distribution assumption of $X\mid Y$ and the independence between $X$ and $({\mathbb{X}},{\mathbb{Y}})$, we have the following control of the estimated discriminant function $\widehat{G}_{\ell}(X)$ uniformly over $\ell\in[L]$.
Theorem 4.
Under the conditions of Theorem 3, we have
$\mathbb{P}\left\{\max_{\ell\in[L]}|\widehat{G}_{\ell}(X)-G_{\ell}(X)|~{}\gtrsim~{}\Delta_{\infty}^{3}\sqrt{sL\log(n\vee p)\log(n)\over n}\right\}\leq 2n^{-\Delta_{\infty}}+n^{-1}.$
Proof.
See Section A.4. ∎
Armed with Theorem 4, we are ready to analyze the proposed classifier $\widehat{g}$ in (2.12) in terms of its excess risk relative to error rate of the Bayes rule in (1.4). Specifically, for any pair $(X,Y)$ from (1.1) that is independent of $({\mathbb{X}},{\mathbb{Y}})$, the excess risk of $\widehat{g}$ is defined as
$\mathcal{R}(\widehat{g}):=\mathbb{P}\left\{Y\neq{\widehat{g}(X)}\right\}-\mathbb{P}\left\{Y\neq{g^{*}(X)}\right\}.$
Note that we use the identification $\{Y=k\}:=\{Y={\bm{e}}_{k}\}$ in the above definition. In the proof of the following theorem, we show that, for any $t>0$,
$\mathcal{R}(\widehat{g})$ $\leq(L-1)~{}t~{}\max_{k\in[L]}\max_{\ell\in[L]\setminus\{k\}}\mathbb{P}\left\{-t\leq G^{(\ell|k)}(X)\leq 0\mid Y={\bm{e}}_{k}\right\}$
$\quad+(L-1)~{}\mathbb{P}\left\{\max_{\ell\in[L]}\left|\widehat{G}_{\ell}(X)-G_{\ell}(X)\right|>t\right\}$
This fact, in combination with Theorem 4, yields the following result on the rates of convergence of $\mathcal{R}(\widehat{g})$.
Theorem 5.
Under the same conditions of Theorem 3 and $\Delta_{\infty}\asymp 1$, we have
$\mathcal{R}(\widehat{g})~{}\lesssim~{}{sL^{2}\log(n\vee p)\log(n)\over n}.$
Proof.
See Section A.5. ∎
Our results allow both the sparsity level $s$ and the number of classes $L$ to grow with the sample size $n$. The existing literature (Cai and Zhang, 2019; Gaynanova, 2020) that analyze the excess risk only consider the case $L\asymp 1$, while Abramovich and Pensky (2019) does allow $s$, $L$ and $n$ to grow, but studies the risk $\mathbb{P}\{Y\neq g(X)\}$ instead of the excess risk $\mathcal{R}(g)$.
Theorem 5 focuses on the statistically most interesting case $\Delta_{\infty}\asymp 1$. Otherwise, if the pairwise separation between different classes grows to infinity, the classification problem becomes much easier and we have super-fast rates of convergence (see, for instance, Cai and Zhang (2019); Bing and Wegkamp (2023)). Indeed, provided
$\min_{k\neq\ell}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell})\asymp\max_{k\neq\ell}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell}),$
using the arguments in the proof of Bing and Wegkamp (2023, Theorem 12), our analysis can be modified easily to show that
$\mathcal{R}(\widehat{g})~{}\lesssim~{}{sL^{2}\log(n\vee p)\log(n)\over n}e^{-c\Delta_{\infty}}+n^{-a}$
for some arbitrary large constant $a>0$ and some constant $c>0$. The first term on the right in the above rate is exponentially fast in $\Delta_{\infty}\to\infty$.
@@ 4 Simulation study
In our simulation study, we evaluate our proposed procedure in Section 2.2 that we term as Linear Discriminant Regularized Regression (LDRR). We also include the procedure in Section 2.3 (called LDRR-F) that is based on Fisher’s discriminant rule after estimating $B$. For estimating $B$ from (1.14), in addition to the lasso penalty (2.7), we also consider the elastic net penalty, $\text{pen}(B)=\lambda\sum_{\ell=1}^{L}[\alpha\|B_{\ell}\|_{2}+(1-\alpha)\|B_{\ell}\|_{1}]$, for some $\alpha\in(0,1)$ and $\lambda>0$ chosen by cross-validation from the R-package glmnet. We compare our proposed methods with the nearest shrunken centroids classifier (PAMR) of Tibshirani et al. (2002), the shrunken centroids regularized discriminant analysis (RDA) method of Guo et al. (2007) and the $\ell_{1}$-penalized linear discriminant (PenalizedLDA) method of Witten and Tibshirani (2011). These methods are available in R-packages pamr, rda, penalizedLDA, and we use their default methods for selecting tuning parameters. Finally, the oracle procedure that uses the true values of ${\mathbb{M}}$, $\Sigma_{w}$, and $\pi$ in the computation of the discriminant functions $G_{k}(x)$ in (2.1) serves as our benchmark.

### v2

@@ 3.3 Application to estimation of $B$ with entry-wise $\ell_{1}$-regularization
In this section we focus on the case where the columns of $B$ are (potentially) sparse and allow for different sparsity patterns across its columns. Therefore, for some positive integer $s\leq p$, we consider the following parameter space of $B$:
@@
For ease of presentation, we simply use the largest number of non-zero entries among all columns of $B$. However, our analysis can be easily extended to $\|B_{\ell}\|_{0}\leq s_{\ell}$ by allowing different $s_{\ell}$ for $\ell\in[L]$. Our analysis also needs the following mild regularity condition on the matrix ${M}$ of conditional means.
@@
There exists an absolute constant $C<\infty$ such that $\|{M}\|_{\infty}\leq C$.
The following theorem establishes explicit rates of $\omega_{2}$ and $\omega_{1}$ in 2 for the regression matrix estimator $\widehat{B}$ in (1.14) using the lasso-penalty in (2.7).
@@
Under model (1.1) with Assumptions 1, 3 and 4, assume that there exists some sufficiently small constant $c>0$ such that $B\in\Theta_{s}$ for some $1\leq s\leq p$ with $(s\vee L)\log(n\vee p)\leq c~{}n.$ Let the estimator $\widehat{B}$ be obtained from (1.14) with any $\lambda$ satisfying
$\lambda\geq C\sqrt{\|\Sigma_{w}\|_{\infty}+\|{M}\|_{\infty}^{2}}\sqrt{\log(n\vee p)\over nL}$ (3.5)
and some finite, large enough constant $C>0$. Then, with probability at least $1-3n^{-1}$, 2 holds for
$\omega_{2}\asymp\lambda\sqrt{sL},\qquad\omega_{1}\asymp\lambda sL.$
@@
See Section B.6. ∎
Our proof is in fact based on a weaker condition than 3 by replacing $\lambda_{p}(\Sigma_{w})$ with the Restricted Eigenvalue condition (RE) on $\Sigma_{w}$. The two main difficulties in our proofs are (a) to control $\max_{j\in[p]}\|{\bm{e}}_{j}^{\top}{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\infty}$, and (b) to bound from below the restricted eigenvalue of the Gram matrix $\widehat{\Sigma}=n^{-1}{\mathbf{X}}^{\top}{\mathbf{X}}$. For (a), as pointed out in Gaynanova (2020), the difficulty in this model is elevated relative to the existing analysis in sparse linear regression models due to the fact ${\mathbf{Y}}$ is not a linear model in ${\mathbf{X}}$. We establish a sharp control of $\max_{j\in[p]}\|{\bm{e}}_{j}^{\top}{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\infty}$ in Lemma 11. For (b), the existing result in Rudelson and Zhou (2012) is not readily applicable as rows of ${\mathbf{X}}$ are generated from a mixture of Gaussians. In Lemma 12 we establish a uniform bound of $v^{\top}(\widehat{\Sigma}-\Sigma)v$ over $s$-sparse vectors $v$ which, in conjunction with the reduction arguments in Rudelson and Zhou (2012), proves (b). Our analysis in this paper allows the number of classes $L$ to grow with the sample size $n$
Gaynanova (2020) considers a group-lasso regularized regression similar to (1.14) and derives the bound of the $\ell_{2}$-norm $\max_{j\in[p]}\|{\bm{e}}_{j}^{\top}{\mathbf{X}}^{\top}({\mathbf{Y}}^{\prime}-{\mathbf{X}}B^{\prime})\|_{2}$ for some different response matrix ${\mathbf{Y}}^{\prime}$ and different target matrix $B^{\prime}$. Their proof is different and in particular requires that $L$ is fixed, independent of $n$.
@@
As an immediate corollary of 4, Theorem 5 and Theorem 6, we have the following upper bound of the misclassification rate of $\widehat{g}$ based on the entry-wise lasso estimator with $\lambda$ chosen as the rate in (3.5).
Corollary 7.
Under the conditions in Theorem 6, assume that $\Delta_{\infty}\asymp 1$. Then
$\mathbb{E}_{{\bm{D}}}\left[\mathcal{R}(\widehat{g})\right]~{}\lesssim~{}\min\left\{{sL^{2}\over n}\log^{2}(n\vee p),~{}1\right\}.$
Our results allow both the sparsity level $s$ and the number of classes $L$ to grow with the sample size $n$. The existing literature (Cai and Zhang, 2019; Gaynanova, 2020) that analyze the excess risk only consider the case $L\asymp 1$, while Abramovich and Pensky (2019) does allow $s$, $L$ and $n$ to grow, but studies the risk $\mathbb{P}\{Y\neq g(X)\}$ instead of the excess risk $\mathcal{R}(g)$. Guarantees on the excess risk are stronger since they together with the Bayes error, which is oftentimes easy to calculate, imply guarantees of the risk.
7 focuses on the statistically most interesting case $\Delta_{\infty}\asymp 1$. Otherwise, if the pairwise separation between different classes grows to infinity, the classification problem becomes much easier and we have super-fast rates of convergence (see, for instance, Cai and Zhang (2019); Bing and Wegkamp (2023)). Indeed, provided $\min_{k\neq\ell}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell})\asymp\max_{k\neq\ell}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell}),$ using the arguments in the proof of Bing and Wegkamp (2023, Theorem 12), our analysis can be modified easily to show that
$\mathbb{E}_{{\bm{D}}}\left[\mathcal{R}(\widehat{g})\right]~{}\lesssim~{}{sL^{2}\over n}\log^{2}(n\vee p)e^{-c\Delta_{\infty}}+n^{-C_{0}}$
for some arbitrary large constant $C_{0}>0$ and some constant $c>0$. The first term on the right in the above rate is exponentially fast in $\Delta_{\infty}\to\infty$.
3.4 Application to estimation of $B$ via reduced-rank regression
We provide analysis for the classifier that estimates the regression coefficient matrix $B$ by reduced-rank regression. Specifically, let $\widehat{B}$ be obtained from (1.14) with $\text{pen}(B)=\lambda~{}{\rm rank}(B)$ for some tuning parameter $\lambda>0$. The following theorem establishes the rates of $\omega_{2}$ and $\omega_{1}$ for which 2 holds. Write $r={\rm rank}(B)$.
@@ 3.4 Application to estimation of $B$ via reduced-rank regression
Under model (1.1) with Assumptions 1 and 3, assume $(p+L)\log(n)\leq c~{}n$ for some sufficiently small constant $c>0$. For the estimator $\widehat{B}$ in (1.14) with any
$\lambda\geq C\sqrt{\|\Sigma_{w}\|_{\rm op}(1+\Delta_{\infty})}\sqrt{(p+L)\log(n)\over nL}$ (3.6)
and some finite, large enough constant $C>0$, with probability at least $1-2n^{-1}$, 2 holds for
$\omega_{2}\asymp\sqrt{\lambda r},\qquad\omega_{1}\asymp\sqrt{\lambda rp}.$
@@
The proof is deferred to Section B.7. ∎
For the reduced-rank estimator, establishing its in-sample prediction risk can be done without any condition on the design matrix (see, for instance, Bunea et al. (2011); Giraud (2011)). However, since 2 is also related with the out-of-sample prediction risk, we need the smallest eigenvalue of the Gram matrix $\widehat{\Sigma}$ to be bounded away from zero. The inequality $(p+L)\log(n)\leq cn$ is assumed for this purpose. In case $p$ is large, it is recommended that an additional ridge penalty be added in $\text{pen}(B)$ to alleviate the singularity issue of $\widehat{\Sigma}$ (Mukherjee and Zhu, 2011). Another main ingredient of our proof of Theorem 8 is to control $\|{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\rm op}$ in Lemma 13. It requires non-standard analysis for the same reason mentioned before that ${\mathbf{Y}}$ is not linearly related with ${\mathbf{X}}$.
Combining 4, Theorem 5 and Theorem 8 immediately yields the following guarantee on the excess risk of the classifier using the reduced-rank estimator of $B$.
Corollary 9.
Under the conditions in Theorem 8, assume $\Delta_{\infty}\asymp 1$ and $\|\Sigma_{w}\|_{\rm op}\lesssim 1$. Then
$\mathbb{E}_{{\bm{D}}}\left[\mathcal{R}(\widehat{g})\right]~{}\lesssim~{}\min\left\{{(p+L)rL\over n}\log^{2}(n),~{}1\right\}.$
Note that for the $p\times L$ matrix $B$ with rank $r$, its effective number of parameters is exactly $(p+L)r$ (Izenman, 1975) which could be much smaller than $pL$ when $r\ll(p\wedge L)$. 9 suggests that the classifier using the reduced-rank estimator has benefit when $r$ is relatively small comparing to $p$ and $L$.
The same remark as the one at the end of Section 3.3 applies, and our analysis can be extended to $\Delta_{\infty}\to\infty$, where the rate in 9 becomes exponentially fast in $\Delta_{\infty}$.
@@ 4 Simulation study
In our simulation study, we evaluate our proposed procedure in Section 2.2 that we term as Linear Discriminant Regularized Regression (LDRR). We also include the procedure in Section 2.3 that is based on Fisher’s discriminant rule after estimating $B$ (called LDRR-F). For estimating $B$ from (1.14), in the sparse scenarios of Section 4.1, we consider both the lasso penalty (2.7) (L1) and the elastic net penalty (L1+L2), $\text{pen}(B)=\lambda\sum_{\ell=1}^{L}[\alpha\|B_{\ell}\|_{1}+(1-\alpha)/2\|B_{\ell}\|_{2}^{2}]$, for some $\alpha\in[0,1]$ and $\lambda>0$ chosen by cross-validation from the R-package glmnet. In the low-rank scenarios of Section 4.2, we consider both the reduced-rank penalty (2.8) (RR) and the reduced-rank penalty plus the ridge penalty (RR+L2), $\text{pen}(B)=\lambda[\alpha~{}{\rm rank}(B)+(1-\alpha)/2\|B\|_{F}^{2}]$, with tuning parameters $\alpha\in[0,1]$ and $\lambda>0$ chosen by cross-validation.
We compare our proposed methods with the nearest shrunken centroids classifier (PAMR) of Tibshirani et al. (2002), the shrunken centroids regularized discriminant analysis (RDA) method of Guo et al. (2007) and the $\ell_{1}$-penalized linear discriminant (PenalizedLDA) method of Witten and Tibshirani (2011). These methods are available in R-packages pamr, rda, penalizedLDA, and we use their default methods for selecting tuning parameters. Finally, the oracle procedure that uses the true values of ${M}$, $\Sigma_{w}$, and $\pi$ in computing the discriminant functions $G_{k}(x)$ in (2.1) serves as our benchmark.
4.1 Sparse scenarios